# MarS-FM: Generative Modeling of Molecular Dynamics via Markov State Models

### v2

@@ 6 Conclusions
Related works. Beyond MD-Emus, our work is related to the class of Boltzmann Generators (BGs), normalizing flows that are trained via the potential energy (Noé et al., 2019; Wirnsberger et al., 2020; Köhler et al., 2021; Rizzi et al., 2021; Garcia Satorras et al., 2021; Rizzi et al., 2023; Midgley et al., 2023; Klein & Noe, 2024; Tan et al., 2025). However, BGs have shown limited scalability beyond small peptides. Alternatively, Jing et al. (2024a); Lewis et al. (2024) proposed sequence-to-structure generative flows trained on MD data, after ignoring any temporal ordering. These methods though suffer from the same data inbalance issues affecting MD-Emus. Learning to sample diverse protein conformations was also investigated in (Jing et al., 2023; Zheng et al., 2024; Lu et al., 2024) and is related to works that perturbed AlphaFold at the MSA level (Del Alamo et al., 2022; Wayment-Steele et al., 2024) or refined its predictions via experiments (Maddipatla et al., 2025). However, these works may fail to attain quantitative distributional matching. Finally, MSMs have been used, for a single peptide dataset, to reweigh samples when fine-tuning a generative model in Lewis et al. (2024).
@@ A.1 Architecture
Inspired by MDGen (Jing et al., 2024b), the velocity network $v_{\theta}:(\mathbb{R}^{21})^{L}\to(\mathbb{R}^{21})^{\text{L}}$, with $L$ being the overall number of residues, employs modified DiT blocks (Jing et al., 2024b; Peebles & Xie, 2023) and Invariant Point Attention (IPA) layers (Jumper et al., 2021). A pseudocode is given in Algorithm 1.
@@
A.2 Drawing trainings pairs
@@ B.1 Datasets
We use the tetrapeptides dataset introduced in Jing et al. (2024b), where trajectories are trained at a resolution of 10 ps over a total duration of 100 ns. We adopt the exact same split as introduced by the authors, comprising 3,109 tetrapeptides for training, 100 for validation, and 100 for testing.
@@
We use the MD-Cath dataset introduced in Mirarchi et al. (2024), which provides molecular dynamics simulations for 5,398 protein domains. Frames are trained at a resolution of 1 ns, with an average trajectory length of 464 ns per replica and the majority of trajectories being 500 ns long. Each domain (ranging from 50 to 500 residues, average 137) is simulated at five temperatures from 320 K to 450 K, with five replicas per temperature, yielding 25 trajectories per domain. For the train/validation/test split, we follow the similar strategy to Lewis et al. (2024) (but more strict) and use MMseqs2 (Steinegger & Söding, 2017) with default parameters and maximum sensitivity to exclude from the test set any domain that shares more than 20% sequence identity with any training or validation domain. This results in a final split of 4,304 domains for training, 538 for validation, and 495 for testing. We primarily evaluate our models on the highest-temperature subset (450 K) of MD-Cath, where most proteins undergo (partial) unfolding within 100–500 ns. In contrast, lower temperature simulations (e.g., at 320 K) often remain near the native structure and exhibit limited conformational diversity—for a quantitative comparison see Appendix C. We use all available replicas for the given temperature.
@@ B.2 Additional details on Markov State Model construction
We follow the MSM construction as described in Section 5. All MSMs are implemented using the deeptime library (Hoffmann et al., 2021). The resulting transition probability matrices are symmetrized post hoc as $\mathbf{P}\leftarrow(\mathbf{P}+\mathbf{P}^{\top})/2$.
@@ B.3.1 Tetrapeptides
We follow the evaluation protocol of Jing et al. (2024b) and compute the Jensen-Shannon Divergence (JSD) between model-generated and reference MD trajectories across several observables: (i) distributions torsional angles, (ii) TICAs, and (iii) equilibrium distributions derived from Markov State Models.
@@
Here, $G_{i}$ denotes the free energy of macrostate $i$, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and $\pi_{i}$ is the stationary distribution (normalized histogram count) of macrostate ii. We use a small floor value (e.g., $10^{-4}$) to avoid numerical instability in the logarithm. Since we do not have access to additional replicas, we did not report the mMAE of the Oracle.
@@ B.3.2 MD-Cath
We follow the exact procedure of Jing et al. (2024a) to compute pairwise Root Mean Square Deviation (RMSD), global Root Mean Square Fluctuation (RMSF), and per-target RMSF. These metrics capture the flexibility and structural variability of conformational ensembles.
@@
  • The Jensen–Shannon Divergence is computed as the squared JSD distance — scipy’s jensenshannon function (Virtanen et al., 2020).
@@
We compute the fraction of residues in secondary structure per frame based on DSSP assignments. Specifically, we include canonical $\alpha$-helix and $\beta$-strand states, using mdtraj’s compute_dssp function (McGibbon et al., 2015),
@@
To estimate folding free energies, we follow the BioEmu protocol (Lewis et al., 2024). For each protein, we first compute a native–contact (FNC) score using all heavy–atom pairs with sequence separation |ij|>3|i-j|>3 and reference distance $d_{ij}^{\text{ref}}<10$ Å. For a given trajectory frame at time $t$, each pair contributes
@@ B.4 Training Details
We follow the hyper-parameters of Jing et al. (2024b) as closely as possible. Throughout all experiments we use an exponential moving average with decay $0.999$, and AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of $1\times 10^{-4}$. We use an architecture described in Appendix A.1 with 5 transformer layers, $384$-dimensional token embeddings, $16$-head multi-head attention, and an IPA layers with 44 heads, $32$-dimensional head size, and 88 query–key as well as 88 value points.
For the tetrapeptide dataset we recreate the setup of Jing et al. (2024b), sampling MDGen-1000 from the released checkpoint and training our MarS-FM model for $1000$ epochs and MDGen-200 for $2500$ epochs, both with a batch size of 88.
On MD-Cath dataset, MarS-FM is run for $1000$ epochs with batch size 88 while sampling training pairs from two source clusters ($x_{0}$) and two corresponding destination clusters ($x_{1}$) with 12 sequences per cluster; MDGen-100 is trained for the same epoch budget with batch size 11, and MDGen-20 for $2500$ epochs with batch size 44.
@@ B.5 Sampling Details
All trajectories are sampled from the first MD frame in the test set; for the MD-Cath data, we use the first frame of the first replica. MDGen is then applied autoregressively, where the last of generated frames at each rollout serves as the conditional input for the next roullout. For MDGen in parallel mode, all calls are conditioned on the initial MD frame; we simply call it multiple times independently.
@@ B.7 Baseline Caveats
For larger systems, we also benchmark against BioEmu (Lewis et al., 2024). Several caveats should be noted when interpreting these results. First, BioEmu uses a backbone-only protein representation, whereas MarS-FM explicitly models both backbone and side-chain torsion angles. Second, BioEmu was trained on substantially more data, including proprietary MD simulations, which makes direct comparisons less controlled. Third, BioEmu currently provides only an inference pipeline, with no publicly available training code. Despite these limitations, we performed inference evaluations on MD-Cath using the pre-trained BioEmu model, generating 500 conformations for comparison with reference MD simulations.
@@ C.1 Supplementary Tetrapeptides TICA plots
C.2 Supplementary plots of generated samples from MD-Cath domains
@@
C.3 JSD for pairwise RMSD and global RMSF, and Pearson $r$ for Folding Free Energies

### v3

@@ 6 Conclusions
Related works. Beyond MD-Emus, our work is related to the class of Boltzmann Generators (BGs), normalizing flows that are trained via the potential energy (noe2019boltzmann; wirnsberger2020targeted; kohler2021smooth; rizzi2021targeted; garcia2021n; rizzi2023multimap; midgley2023se; kleintransferable; tan2025scalable). However, BGs have shown limited scalability beyond small peptides. Alternatively, jing2024alphafold; lewis2024scalable proposed sequence-to-structure generative flows trained on MD data, after ignoring any temporal ordering. These methods though suffer from the same data inbalance issues affecting MD-Emus. Learning to sample diverse protein conformations was also investigated in (jing2023eigenfold; zheng2024predicting; lu2024structure) and is related to works that perturbed AlphaFold at the MSA level (del2022sampling; wayment2024predicting) or refined its predictions via experiments (maddipatla2025inverse). However, these works may fail to attain quantitative distributional matching. Finally, MSMs have been used, for a single peptide dataset, to reweigh samples when fine-tuning a generative model in lewis2024scalable.
@@ A.1 Architecture
Inspired by MDGen (jing24generative), the velocity network $v_{\theta}:(\mathbb{R}^{21})^{L}\to(\mathbb{R}^{21})^{\text{L}}$, with $L$ being the overall number of residues, employs modified DiT blocks (jing24generative; peebles2023scalable) and Invariant Point Attention (IPA) layers (jumper2021highly). A pseudocode is given in Algorithm 1.
@@
A.2 Drawing training pairs
@@ B.1 Datasets
We use the tetrapeptides dataset introduced in jing24generative, where trajectories are trained at a resolution of 10 ps over a total duration of 100 ns. We adopt the exact same split as introduced by the authors, comprising 3,109 tetrapeptides for training, 100 for validation, and 100 for testing.
@@
We use the MD-Cath dataset introduced in mirarchi2024mdcath, which provides molecular dynamics simulations for 5,398 protein domains. Frames are trained at a resolution of 1 ns, with an average trajectory length of 464 ns per replica and the majority of trajectories being 500 ns long. Each domain (ranging from 50 to 500 residues, average 137) is simulated at five temperatures from 320 K to 450 K, with five replicas per temperature, yielding 25 trajectories per domain. For the train/validation/test split, we follow the similar strategy to lewis2024scalable (but more strict) and use MMseqs2 (steinegger2017mmseqs2) with default parameters and maximum sensitivity to exclude from the test set any domain that shares more than 20% sequence identity with any training or validation domain. This results in a final split of 4,304 domains for training, 538 for validation, and 495 for testing. We primarily evaluate our models on the highest-temperature subset (450 K) of MD-Cath, where most proteins undergo (partial) unfolding within 100–500 ns. In contrast, lower temperature simulations (e.g., at 320 K) often remain near the native structure and exhibit limited conformational diversity—for a quantitative comparison see Appendix C. We use all available replicas for the given temperature.
@@ B.2 Additional details on Markov State Model construction
We follow the MSM construction as described in Section 5. All MSMs are implemented using the deeptime library (hoffmann2021deeptime). The resulting transition probability matrices are symmetrized post hoc as $\mathbf{P}\leftarrow(\mathbf{P}+\mathbf{P}^{\top})/2$.
@@ B.3.1 Tetrapeptides
We follow the evaluation protocol of jing24generative and compute the Jensen-Shannon Divergence (JSD) between model-generated and reference MD trajectories across several observables: (i) distributions torsional angles, (ii) TICAs, and (iii) equilibrium distributions derived from Markov State Models.
@@
Here, $G_{i}$ denotes the free energy of macrostate $i$, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, and $\pi_{i}$ is the stationary distribution (normalized histogram count) of macrostate $i$. We use a small floor value (e.g., $10^{-4}$) to avoid numerical instability in the logarithm. Since we do not have access to additional replicas, we did not report the mMAE of the Oracle.
@@ B.3.2 MD-Cath
We follow the exact procedure of jing2024alphafold to compute pairwise Root Mean Square Deviation (RMSD), global Root Mean Square Fluctuation (RMSF), and per-target RMSF. These metrics capture the flexibility and structural variability of conformational ensembles.
@@
  • The Jensen–Shannon Divergence is computed as the squared JSD distance — scipy’s jensenshannon function (2020SciPy-NMeth).
@@
We compute the fraction of residues in secondary structure per frame based on DSSP assignments. Specifically, we include canonical $\alpha$-helix and $\beta$-strand states, using mdtraj’s compute_dssp function (McGibbon2015MDTraj),
@@
To estimate folding free energies, we follow the BioEmu protocol (lewis2024scalable). For each protein, we first compute a native–contact (FNC) score using all heavy–atom pairs with sequence separation |ij|>3|i-j|>3 and reference distance $d_{ij}^{\text{ref}}<10$ Å. For a given trajectory frame at time $t$, each pair contributes
@@ B.4 Training Details
We follow the hyper-parameters of jing24generative as closely as possible. Throughout all experiments we use an exponential moving average with decay $0.999$, and Adam (kingma2014adam) optimizer with a learning rate of $1\times 10^{-4}$. We use an architecture described in Appendix A.1 with 5 transformer layers, $384$-dimensional token embeddings, $16$-head multi-head attention, and an IPA layers with $4$ heads, $32$-dimensional head size, and $8$ query–key as well as $8$ value points.
For the tetrapeptide dataset we recreate the setup of jing24generative, sampling MDGen-1000 from the released checkpoint and training our MarS-FM model for $1000$ epochs and MDGen-200 for $2500$ epochs, both with a batch size of $8$.
On MD-Cath dataset, MarS-FM is run for $1000$ epochs with batch size $8$ while sampling training pairs from two source clusters ($x_{0}$) and two corresponding destination clusters ($x_{1}$) with 12 sequences per cluster; MDGen-100 is trained for the same epoch budget with batch size $1$, and MDGen-20 for $2500$ epochs with batch size $4$.
At inference, we generate samples by integrating the learned velocity field using the adaptive dopri5 solver via torchdiffeq (chen2018neural), with absolute tolerance $10^{-6}$, relative tolerance $10^{-3}$, and 50 integration steps.
@@ B.5 Sampling Details
All trajectories are sampled from the first MD frame in the test set; for the MD-Cath data, we use the first frame of the first replica. MDGen is then applied autoregressively, where the last of generated frames at each rollout serves as the conditional input for the next rollout. For MDGen in parallel mode, all calls are conditioned on the initial MD frame; we simply call it multiple times independently.
@@ B.7 Baseline Caveats
For larger systems, we also benchmark against BioEmu (lewis2024scalable). Several caveats should be noted when interpreting these results. First, BioEmu uses a backbone-only protein representation, whereas MarS-FM explicitly models both backbone and side-chain torsion angles. Second, BioEmu was trained on substantially more data, including proprietary MD simulations, which makes direct comparisons less controlled. Third, BioEmu currently provides only an inference pipeline, with no publicly available training code. Despite these limitations, we performed inference evaluations on MD-Cath using the pre-trained BioEmu model, generating 500 conformations for comparison with reference MD simulations.
@@ C.1 Supplementary Tetrapeptides TICA plots
C.2 Additional dynamical observables for tetrapeptides
To complement the thermodynamic metrics in the main text, we report additional dynamical observables on the 4AA tetrapeptide benchmark. Following MDGen (jing24generative), we consider the normalized torsion autocorrelation
$C_{\theta}(\Delta t)\;=\;\big\langle\cos\big(\theta_{t}-\theta_{t+\Delta t}\big)\big\rangle,$
and define a relaxation time $\tau_{\theta}$ as the lag at which $C_{\theta}(\Delta t)$ decays below $1/e$ (up to $10\,\text{ns}$).
Figure 8 compares backbone and sidechain decorrelation curves for representative test peptides (ALDA, IWHF, KSIY, PGKM, SSNN) across MDGen-1000, MARS-FM, and the hybrid MarS-FM $\oplus$ MDGen-200 scheme. Figure 7 correlates model and MD relaxation times for all backbone and sidechain torsions in the test set.
[…]
MDGen-1000
MarS-FM
[…]
Figure 7: Torsion relaxation times for 4AA tetrapeptides. Scatter plots of model versus MD relaxation times $\tau_{\theta}$ for backbone (blue) and sidechain (orange) torsions. Axes are logarithmic; the dashed line indicates $y=x$. Pearson correlations $r$ for backbone and sidechain torsions are reported in the legend of each panel.
[…]
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 8: Backbone and sidechain torsion decorrelation for 4AA tetrapeptides. Normalized autocorrelation $C_{\theta}(\Delta t)$ as a function of lag time for backbone (left panels in each image) and sidechain (right panels) torsions in five representative test peptides (rows: ALDA, IWHF, KSIY, PGKM, SSNN) and three models (columns: MDGen-1000, $MarS-FM \oplusMDGen-$200). Solid lines denote MD reference trajectories, dashed lines the corresponding model trajectories.
C.3 Supplementary plots of generated samples from MD-Cath domains
@@
C.4 Ablation on hierarchical tree sampling
MarS-FM samples MD-Cath trajectories using hierarchical tree sampling (§5.2), where each node spawns $N$ children in parallel before the tree is further expanded. To study the impact of this branching factor, we fix the total sample budget to 500 conformations per domain on the 450 K MD-Cath test set (§5.2) and vary $N\in\{10,20,50,100,250,500\}$. We then evaluate the same structural and thermodynamic metrics as in Table 2.
Results are reported in Table 5. Very small trees ($N=10$) require more expansion depth to reach the target sample budget, increasing the number of autoregressive steps and leading to noticeable degradation in structural flexibility and folding free energies. Increasing $N$ to intermediate values ($N\approx 100$$250$) consistently improves RMSD/RMSF correlations and reduces divergences in radius of gyration and secondary-structure distributions. However, pushing $N$ too high ($N=500$) makes the tree very shallow: most samples are generated close to the initial frame, which harms state-space exploration and MSM reconstruction, as reflected by the higher MSM JSD and weaker RMSF correlations. Overall, intermediate branching factors around N=100N=100250250 provide the best trade-off between limiting temporal error accumulation and accurately recovering the MSM and long-timescale observables; our default choice of $N=200$ used in the main experiments lies in this regime.
Table 5: Effect of branching factor NN in hierarchical MarS-FM tree sampling on the MD-Cath 450 K test set. We report Pearson $r\uparrow$ for RMSD and RMSF, forward KL divergence $\downarrow$, JSD $\downarrow$ of gyration radius, secondary-structure fractions, and MSM distributions, and folding free energy MAE (kcal/mol) $\downarrow$. Results are based on 500 generated conformations per domain and averaged over 5 inference runs.
Pairwise RMSD $r$ Global RMSF $r$ Per target RMSF $r$ Gyration Radius KL Gyration Radius JSD Secondary Structure KL Secondary Structure JSD MSM JSD $\Delta G_{\text{fold}}$ MAE $N=10$ $0.54$ ± $\pm\,.007$ $0.64$ ± $\pm\,.005$ $0.89$ ± $\pm\,.002$ $0.56$ ± $\pm\,.001$ $0.13$ ± $\pm\,.001$ $2.12$ ± $\pm\,.018$ $0.24$ ± $\pm\,.001$ $0.17$ ± $\pm\,.002$ $2.38$ ± $\pm\,.005$ $N=20$ $0.62$ ± $\pm\,.005$ $0.69$ ± $\pm\,.005$ $0.90$ ± $\pm\,.002$ $0.51$ ± $\pm\,.002$ $0.11$ ± $\pm\,.000$ $1.54$ ± $\pm\,.003$ $0.18$ ± $\pm\,.001$ $0.18$ ± $\pm\,.007$ $1.73$ ± $\pm\,.010$ $N=50$ $0.65$ ± $\pm\,.001$ $0.71$ ± $\pm\,.001$ $0.90$ ± $\pm\,.002$ $0.52$ ± $\pm\,.003$ $0.10$ ± $\pm\,.000$ $1.31$ ± $\pm\,.007$ $0.17$ ± $\pm\,.000$ $0.17$ ± $\pm\,.003$ $1.43$ ± $\pm\,.019$ $N=100$ $0.65$ ± $\pm\,.002$ $0.71$ ± $\pm\,.002$ $0.90$ ± $\pm\,.003$ $0.52$ ± $\pm\,.002$ $0.10$ ± $\pm\,.001$ $1.13$ ± $\pm\,.007$ $0.15$ ± $\pm\,.001$ $0.18$ ± $\pm\,.002$ $1.23$ ± $\pm\,.016$ $N=250$ $0.65$ ± $\pm\,.003$ $0.71$ ± $\pm\,.002$ $0.89$ ± $\pm\,.002$ $0.57$ ± $\pm\,.003$ $0.10$ ± $\pm\,.001$ $0.87$ ± $\pm\,.007$ $0.12$ ± $\pm\,.001$ $0.19$ ± $\pm\,.005$ $1.03$ ± $\pm\,.003$ $N=500$ $0.61$ ± $\pm\,.001$ $0.69$ ± $\pm\,.002$ $0.84$ ± $\pm\,.003$ $0.87$ ± $\pm\,.003$ $0.12$ ± $\pm\,.000$ $0.83$ ± $\pm\,.010$ $0.12$ ± $\pm\,.001$ $0.25$ ± $\pm\,.004$ $1.20$ ± $\pm\,.004$
C.5 Ablation on MSM state construction for MD-Cath 450 K
In the main MD-Cath experiments (§5.2), we construct per-domain MSMs by $k$-means clustering of normalized radius of gyration and secondary-structure fractions into $c=10$ metastable states with lag time $l=50$ ns. To assess the sensitivity of MarS-FM to this choice, we perform an ablation over the MSM discretization on the 450 K test set, varying both the number of clusters $c$ and the MSM lag time $l$ while keeping all other settings fixed.
Specifically, we retrain MarS-FM and the hybrid MarS-FM $\oplus$ MDGen-20 using MSMs built with $(c,l)\in\{(10,50),(10,100),(10,20),(20,50),(5,50)\}$, where $c$ denotes the number of $k$-means clusters and $l$ the lag time (in ns) used to estimate the transition matrix. For each configuration we generate 500 conformations per test domain. Results are reported in Table 6.
Overall, we observe that MarS-FM is robust to reasonable changes in the MSM discretization. For both MarS-FM and MarS-FM $\oplus$ MDGen-20, all $(c,l)$ settings yield very similar RMSD/RMSF correlations and divergences on radius of gyration and secondary-structure distributions. Increasing the MSM lag time from 50 ns to 100 ns slightly improves MSM JSD and $\Delta G_{\text{fold}}$ MAE, while shortening the lag to 20 ns degrades the gyration-radius KL and MSM JSD for both models. Varying the number of clusters at fixed lag time has only modest effects: coarse discretizations with $c=5$ tend to worsen secondary-structure KL/JSD and MSM JSD, whereas $c=20$ slightly improves MSM recovery without changing the overall conclusions. Taken together, these results indicate that the default MSM choice $(c=10,l=50\penalty 10000\ \text{ns})$ used in the main text offers a good trade-off across metrics, and that our improvements over MD-Emus do not rely on a finely tuned state construction.
Table 6: Ablation on MSM discretization for MD-Cath 450 K. Pearson $r\uparrow$ for RMSD and RMSF, forward KL divergence $\downarrow$ and JSD $\downarrow$ of gyration radius, fraction of secondary structures, MSMs distribution, and folding free energy MAE (kcal/mol) $\downarrow$. Results based on sampled trajectories of 500 conformations compared to ground-truth distributions. MSMs are labeled by the number of clusters $c$ and lag time (e.g., $c=10$, lag 50 ns).
MSM (clusters, lag) Model Pairwise RMSD r Global RMSF r Per target RMSF r Gyration Radius KL Gyration Radius JSD Secondary Structures KL Secondary Structures JSD MSM JSD $\Delta G_{\text{fold}}$ MAE $c=10$, lag 50 ns MarS-FM $\oplus$ MDGen-20 0.63 0.69 0.83 0.98 0.13 0.73 0.11 0.24 1.02 MarS-FM 0.65 0.71 0.89 0.55 0.10 0.93 0.13 0.19 1.05 $c=10$, lag 100 ns MarS-FM $\oplus$ MDGen-20 0.69 0.73 0.85 0.70 0.12 0.74 0.12 0.20 0.91 MarS-FM 0.69 0.73 0.90 0.50 0.11 1.05 0.15 0.15 1.21 $c=10$, lag 20 ns MarS-FM $\oplus$ MDGen-20 0.56 0.66 0.82 1.07 0.14 0.75 0.12 0.25 0.97 MarS-FM 0.53 0.65 0.86 0.72 0.11 1.17 0.16 0.21 0.98 $c=20$, lag 50 ns MarS-FM $\oplus$ MDGen-20 0.64 0.69 0.83 0.94 0.13 0.76 0.12 0.24 0.93 MarS-FM 0.59 0.67 0.89 0.64 0.11 1.10 0.14 0.20 0.96 $c=5$, lag 50 ns MarS-FM $\oplus$ MDGen-20 0.62 0.68 0.83 0.88 0.13 0.93 0.14 0.24 0.95 MarS-FM 0.57 0.65 0.89 0.57 0.10 1.58 0.18 0.20 1.23
@@ C.5 Ablation on MSM state construction for MD-Cath 450 K
C.6 JSD for pairwise RMSD and global RMSF, and Pearson $r$ for Folding Free Energies